Solve this system of equation using substitution, 2x+y=3 x+2y=negetive3
Answer by checkley77(12844)

2x+y=3
x+2y=-3
x=-2y-3
Now replace x in the first equation by (-2y-3).
2(-2y-3)+y=3
-4y-6+y=3
-3y=3+6
-3y=9
y=9/-3
y=-3 ans.
2x-2=3
2x=3=3
2x=6
x=6/2
x=3 ans.